Given that income is $500 and PX = $20 and PY = $5, what is the market rate of substitution between goods X and Y? Show solutions



1
Expert's answer
2022-02-17T11:11:18-0500

Equilbrium condition,

"\frac{MUx}{MUy} =\frac{Px}{Py}"

"=- \frac{20}{5}"

"= -4"

Marginal rate of substitution between two goods is a negative ratio of the prices of these goods. Mathematically, it means that a price of one good is divided by a price of another good, with a negative sign. Similarly, marginal rate of substitution is equal to the slope of the budget constraint line.

That's why there is a negative sign.